OpenStudy (anonymous):

Solve the inequality x 15 --- + ---- <= 4 4 x+7 for domain

OpenStudy (amistre64):

x cant be -7 so we can rule any option that includes that...

OpenStudy (amistre64):

we can get like denoms again; that worked prety well last time :) x^2 +7x + 60 <= 16x +16(17) x^2 -9x +60 - 16(17) <= 0 ..... 142 + 60 + 70 = 272 - 60 = -212 x^2 -9x -212 <=0

OpenStudy (amistre64):

9/2 +- sqrt(81 -4(-212))/2 9/2 +- sqrt(929)/2..oh you picked a doozey eh...

OpenStudy (amistre64):

whats your options?

OpenStudy (amistre64):

x^2 +7x + 60 <= 16x +16(17) x^2 -9x +60 - 16(7) <= 0 ..... 70+42 = 112 - 60 = 52 x^2 -9x -52 <=0 is better 9/2 +- sqrt(81+108)/2 9/2 +- sqrt(189)/2 [ 9 (+-) 3sqrt(21) ]/2